Jane Windsor financed a $5,900 ski boat with a 12% add-on interest installment loan for 12 months. Given the loan required a 10% down payment, determine the following:
The amount of the finance charge?
The amount of the finance charge rebate if the loan were to be paid after the 10th payment?

Answer by Theo(13342)

from a mathematical perspective, i would estimate the following:
the loan is .9 * 5900 = 5310.
the interest on the loan is .12 * 5310 = 637.2
this interest charge is divided by 12 to be equal to 53.1 each month
if the loan is paid off in 10 months, there would be 2 months of interest payments remaining.
the rebate would therefore be 2 * 53.1 = 106.2.

note that an add on loan is a simple interest type loan.
the formula for simple interest type loans is.
i = p * r * t
i is the interest
p is the principal
r is the interest rate per time period
t is the number of time periods

if the number of time periods are in months, then the annual interest rate is divided by 12.

with the full term of the loan, the formula becomes i = 5310 * .01 * 12 = 637.2
with the term of the loan equal to 10 months, the formula becomes i = 5310 * .01 * 10 = 531
the difference is 637.2 minus 531 = 106.2
that's what would get refunded from a mathematical perspective.
i have no idea how the banks would handle this.